the surface area of a cylinder is 180 π square inches and the height is 9 inches. Find the radius.
Answer by jorel1380(2518)

Surface Area=2πr2+2πrh
"=2πr(r+h)
180π=2πr(r+h)
90=r2+9r
r2+9r-90=0
(r+15)(r-6)=0
r=-15 or 6
Throwing out the negative result, we get radius equals 6 in..